The fundamental frequency of a sting stretched with a weight of 4 kg is 256 Hz. The weight required to produce its octave is
(a) 4 kg wt
(b) 12 kg wt
(c) 16 kg wt
(d) 24 kg wt

Grade:Upto college level

1 Answers

Vikas TU
3 years ago
If the fundamental frequency of the spring is f
Then frequency of its octave is 2f.
f = K (T)1/2
2f = K(T')1/2
Hence T' = 4T
T' = 4*4 = 16 kg- wt